# Fast Penalized Generalized Estimating Equations for Large Longitudinal Functional Datasets

### v1

@@ 2.4 Tuning $\Lambda$
where $\tilde{n}_{k}=\frac{\sum_{i=1}^{N}n_{i}}{\sum_{i\not\in\mathcal{K}_{k}}n_{i}}$. By using the full sample estimate $\left[\frac{1}{N}\sum_{i=1}^{N}\mathbb{W}_{i}(\widehat{\bm{\theta}}^{(0)}_{\Lambda_{0}})+\Lambda_{1}\mathbb{S}\right]^{-1}$, we only need to invert this $p\times p$ matrix once for each value of $\Lambda_{1}$, instead of inverting a fold-specific $p\times p$ matrix for each unique $\{k,\Lambda_{1}\}$ pair. The strategy of keeping $\widehat{\bm{\theta}}^{(0)}_{{\Lambda_{0}}}$ and $\left[\frac{1}{N}\sum_{i=1}^{N}\mathbb{W}_{i}(\widehat{\bm{\theta}}^{(0)}_{\Lambda_{0}})+\Lambda_{1}\mathbb{S}\right]^{-1}$ fixed across folds is motivated by an analogous strategy for cluster bootstrapping of unpenalized one-step GEE (see Remark and Theorem 3.3 in Cheng et al., (2013)). Specifically, Cheng et al., (2013) showed that a cluster bootstrap that fixes these two quantities (at the full-sample estimates) across replicates enjoys the same theoretical guarantees asymptotically as an approach that re-estimates these quantities in each replicate-specific sample. In our simulations, our adaptation of this strategy for cluster CV was often dramatically faster than, and performed nearly identically to, a CV strategy that calculates $\widehat{\bm{\theta}}^{k}_{{\Lambda_{1}}}$ using the fold-specific estimate $\left[\frac{1}{N-|\mathcal{K}_{k}|}\sum_{i\not\in\mathcal{K}_{k}}\mathbb{W}_{i}(\widehat{\bm{\theta}}^{(0)}_{\Lambda_{0}})+\left(\frac{\sum_{i\not\in\mathcal{K}_{k}}n_{i}}{\sum_{i=1}^{N}n_{i}}\right)\Lambda_{1}\mathbb{S}\right]^{-1}$. Fourth, we avoid tuning over a large grid of $\Lambda_{1}$ values by using a sequential CV procedure (see Appendix A.1 for details). We found these strategies performed well with $K=10$ in our simulations and data application.
2.5 Coefficient Estimator Variance
2.5.1 Sandwich Variance Estimator
We estimate the variance of our one-step estimator using the sandwich form
$\widehat{\bm{V}}_{\Lambda}(\bm{\theta})=\frac{1}{N}[\widehat{\bm{H}}_{\Lambda}(\bm{\theta})]^{-1}\widehat{\bm{M}}_{\Lambda}(\bm{\theta})[\widehat{\bm{H}}_{\Lambda}(\bm{\theta})]^{-1},$ (6)
where $\widehat{\bm{M}}_{\Lambda}(\bm{\theta})=\mathbb{P}_{N}[\bm{U}_{\Lambda}(\mathbf{X}_{i},\bm{Y}_{i};\bm{\theta})\bm{U}_{\Lambda}(\mathbf{X}_{i},\bm{Y}_{i};\bm{\theta})^{T}]$, and $\widehat{\bm{H}}_{\Lambda}(\bm{\theta})=\mathbb{P}_{N}[\mathbb{D}_{i}(\bm{\theta})^{T}\widehat{\mathbb{V}}_{i}(\bm{\theta})^{-1}\mathbb{D}_{i}(\bm{\theta})]+\Lambda\mathbb{S}$, and $\mathbb{P}_{N}[f(O_{i})]=\frac{1}{N}\sum_{i=1}^{N}f(O_{i})$ denotes the sample average of a given function $f$. We estimate $\widehat{\text{Var}}(\widehat{\bm{\theta}}_{\Lambda_{1}}^{(1)})$ via $\widehat{\bm{V}}_{\Lambda_{1}}(\widehat{\bm{\theta}}_{\Lambda_{1}}^{(1)})$, i.e., by plugging in $\widehat{\bm{\theta}}^{(1)}_{\Lambda_{1}}$, $\widehat{\mathbb{V}}_{i}(\widehat{\bm{\theta}}^{(1)}_{\Lambda_{1}})$, $\widehat{\mathbf{A}}_{i}(\widehat{\bm{\theta}}^{(1)}_{\Lambda_{1}})$, and $\Lambda_{1}$.
2.5.2 Fast Cluster Bootstrap Variance Estimator
@@
Motivated by theory for cluster bootstrapping in non-functional one-step GEE (Cheng et al.,, 2013), we propose a fast cluster bootstrap as an alternative method to estimate $\text{Var}(\widehat{\bm{\theta}}_{{\Lambda_{1}}}^{(1)})$, or to construct non-parametric bootstrap-based joint CIs. Namely, for bootstrap replicate, $t$

### v2

@@ A.6.1 Fast Cluster Cross-Validation
where $\tilde{n}_{k}=\frac{\sum_{i=1}^{N}n_{i}}{\sum_{i\not\in\mathcal{K}_{k}}n_{i}}$. By using the full sample estimate $\left[\frac{1}{N}\sum_{i=1}^{N}\mathbb{W}_{i}(\widehat{\bm{\theta}}^{(0)}_{\Lambda_{0}})+\Lambda_{1}\mathbb{S}\right]^{-1}$, we only need to invert this $p\times p$ matrix once for each value of $\Lambda_{1}$, instead of inverting a fold-specific $p\times p$ matrix for each unique $\{k,\Lambda_{1}\}$ pair. The strategy of keeping $\widehat{\bm{\theta}}^{(0)}_{\Lambda_{0}}$ and $\left[\frac{1}{N}\sum_{i=1}^{N}\mathbb{W}_{i}(\widehat{\bm{\theta}}^{(0)}_{\Lambda_{0}})+\Lambda_{1}\mathbb{S}\right]^{-1}$ fixed across folds is motivated by an analogous strategy for cluster bootstrapping of unpenalized one-step GEE (see Remark and Theorem 3.3 in Cheng et al. (2013)). Specifically, Cheng et al. (2013) showed that a cluster bootstrap that fixes these two quantities (at the full-sample estimates) across replicates enjoys the same theoretical guarantees asymptotically as an approach that re-estimates these quantities in each replicate-specific sample. In our simulations, our adaptation of this strategy for cluster CV was often dramatically faster than, and performed nearly identically to, a CV strategy that calculates $\widehat{\bm{\theta}}^{k}_{\Lambda_{1}}$ using the fold-specific estimate $\left[\frac{1}{N-|\mathcal{K}_{k}|}\sum_{i\not\in\mathcal{K}_{k}}\mathbb{W}_{i}(\widehat{\bm{\theta}}^{(0)}_{\Lambda_{0}})+\left(\frac{\sum_{i\not\in\mathcal{K}_{k}}n_{i}}{\sum_{i=1}^{N}n_{i}}\right)\Lambda_{1}\mathbb{S}\right]^{-1}$. Fourth, we avoid tuning over a large grid of $\Lambda_{1}$ values by using a sequential CV procedure (see Appendix A.6.3 below for details). We found these strategies performed well with $K=10$ in our simulations and data application. We describe the fast cluster CV procedure in Algorithm S2.
Algorithm S2 Fast cluster CV for one-step fGEE
1:Inputs:
  • Initial estimate $\widehat{\bm{\theta}}^{(0)}_{\bm{\Lambda}_{0}}$
  • Penalty matrix $\mathbb{S}$
  • Cluster folds $\{\mathcal{K}_{1},\dots,\mathcal{K}_{K}\}$ where for $K\geq 2$ $\mathcal{K}_{k}\subset[N],~\cup_{k}\mathcal{K}_{k}=[N],~\cap_{k}\mathcal{K}_{k}=\emptyset$
  • Candidate grid $\mathcal{G}$ of smoothing parameter matrices $\bm{\Lambda}$
  • Precomputed $\mathbb{W}_{i}(\widehat{\bm{\theta}}^{(0)}_{\Lambda_{0}}),\mathbf{b}_{i}(\widehat{\bm{\theta}}^{(0)}_{\Lambda_{0}})\}_{i=1}^{N}$
  • $\mathrm{FoldLoss}(\cdot)$: cross-validated error criteria (e.g., squared-error, negative log-likelihood)
    • $\mathbb{W}_{i}\equiv\mathbb{W}_{i}(\widehat{\bm{\theta}}^{(0)}_{\Lambda_{0}})=\{\mathbb{D}_{i}(\widehat{\bm{\theta}}^{(0)}_{\Lambda_{0}})\}^{T}\{\widehat{\mathbb{V}}_{i}(\widehat{\bm{\theta}}_{\Lambda_{0}}^{(0)})\}^{-1}\{\mathbb{D}_{i}(\widehat{\bm{\theta}}^{(0)}_{\Lambda_{0}})\}\in\mathbb{R}^{p\times p}$
    • $\mathbf{b}_{i}\equiv\mathbf{b}_{i}(\widehat{\bm{\theta}}^{(0)}_{\Lambda_{0}})=\{\mathbb{D}_{i}(\widehat{\bm{\theta}}^{(0)}_{\Lambda_{0}})\}^{T}\{\widehat{\mathbb{V}}_{i}(\widehat{\bm{\theta}}_{\Lambda_{0}}^{(0)})\}^{-1}\left\{\bm{Y}_{i}-\widehat{\bm{\mu}}_{i}(\widehat{\bm{\theta}}^{(0)}_{\Lambda_{0}})\right\}\in\mathbb{R}^{p}$
2:for each $\bm{\Lambda}\in\mathcal{G}$ do
3:  $\bm{H}_{\mathrm{glob}}(\bm{\Lambda})\leftarrow\Big(\frac{1}{N}\sum_{i=1}^{N}\mathbb{W}_{i}\Big)+\bm{\Lambda}\,\mathbb{S}$.
4:  Invert $\bm{H}_{\mathrm{glob}}(\bm{\Lambda})$ and store $\bm{H}_{\mathrm{glob}}(\bm{\Lambda})^{-1}$
5:  for $k=1$ to $K$ do
6:    $\tilde{n}_{k}\leftarrow\Big(\sum_{i=1}^{N}n_{i}\Big)\Big/\Big(\sum_{i\notin\mathcal{K}_{k}}n_{i}\Big)$
7:    $\widehat{\bm{\theta}}^{(1)}_{k}(\bm{\Lambda})\leftarrow\widehat{\bm{\theta}}^{(0)}_{\bm{\Lambda}_{0}}+\bm{H}_{\mathrm{glob}}(\bm{\Lambda})^{-1}\Big[\frac{1}{N}\sum_{i\notin\mathcal{K}_{k}}\big\{\tilde{n}_{k}\,\mathbf{b}_{i}-\bm{\Lambda}\,\mathbb{S}\,\widehat{\bm{\theta}}^{(0)}_{\bm{\Lambda}_{0}}\big\}\Big$]
8:    $\mathrm{NLL}_{k}(\bm{\Lambda})\leftarrow\mathrm{FoldLoss}\big(\widehat{\bm{\theta}}^{(1)}_{k}(\bm{\Lambda}),\,\mathcal{K}_{k}\big)$
9:  $\mathrm{CV}(\bm{\Lambda})\leftarrow\frac{1}{K}\sum_{k=1}^{K}\mathrm{NLL}_{k}(\bm{\Lambda})$
[…]
11:return $\bm{\Lambda}^{\star}$
A.6.2 Full-Sample Cross Validation
We define a standard cluster (K-Fold) cross-validation one-step estimator as
where ${n}^{*}_{k}=\frac{\sum_{i\not\in\mathcal{K}_{k}}n_{i}}{\sum_{i=1}^{N}n_{i}}$. In preliminary simulations (not shown here), one-step fGEE estimates fit with smoothing parameters selected with the full-sample and fast cluster CVs produced nearly identical functional coefficient RMSE and empirical 95% CI coverage values. For that reason, all results shown in the main text and Appendix are obtained from fGEEs fit with smoothing parameters selected with the fast cluster CV.
A.6.3 Sequential Tuning Procedure
The diagonal smoothing matrix $\Lambda$ contains smoothing parameters $\lambda_{1},\ldots,\lambda_{q}$, each repeated based on the number of knots used for its functional coefficient. To identify the correct range for the $\lambda_{1},\ldots,\lambda_{q}$, we apply an iterative CV strategy that is designed to be fast as $q$ grows. By dividing the tuning into three stages, we avoid tuning over a large $q$ dimensional grid that can be computationally impractical even for $q\geq 3$. In step (1), we tune over a small one-dimensional grid to identify the correct order of magnitude for the smoothing parameters: $\bm{\Lambda}_{(1)}=\{\alpha_{1}\Lambda^{(0)},\ldots,\alpha_{L}\Lambda^{(0)}\}$ where, for example, $\{\alpha_{1},\ldots,\alpha_{L_{1}}\}=\{0.001,0.01,0.1,1,10,100,1000\}$. This exploits the fact that the unique diagonal entries $\lambda_{1}^{(0)},\ldots,\lambda_{q}^{(0)}$ of the $\Lambda_{0}$ selected by REML are, in our experience, on a reasonable relative scale. In step (2), we tune over a small $q$-dimensional grid constructed around the smoothing parameter values that minimize the cross-validated criteria (e.g. MSE, negative log-likelihood) in step (1). Denoting $\tilde{\lambda}^{(1)}_{1},\ldots,\tilde{\lambda}^{(1)}_{q}$ as these selected values, we tune over a grid with all unique combinations (with the R function $\texttt{expand.grid}()$) of grid $\bm{\Lambda}_{(2)}=\left\{\{\alpha_{1}\tilde{\lambda}^{(1)}_{1},\ldots,\alpha_{L_{2}}\tilde{\lambda}^{(1)}_{1}\},\ldots,\{\alpha_{1}\tilde{\lambda}^{(1)}_{q},\ldots,\alpha_{L}\tilde{\lambda}^{(1)}_{q}\}\right\}$, where, for example, $\{\alpha_{1},\ldots,\alpha_{L_{2}}\}=$
$\{0.001,0.01,0.1,1,10,100,1000\}$. Step (2) identifies a reasonable order of magnitude for each smoothing parameter with a small grid. Finally, in step (3), we tune each smoothing parameter within a local neighborhood around the selected values from step (2), $\tilde{\lambda}^{(2)}_{1},\ldots,\tilde{\lambda}^{(2)}_{q}$. Specifically, we tune over the unique combinations, $\bm{\Lambda}_{(3)}=\left\{\{\alpha_{1}\tilde{\lambda}^{(2)}_{1},\ldots,\alpha_{L_{3}}\tilde{\lambda}^{(2)}_{1}\},\ldots,\{\alpha_{1}\tilde{\lambda}^{(2)}_{q},\ldots,\alpha_{L_{3}}\tilde{\lambda}^{(2)}_{q}\}\right\}$, where, for example, $\{\alpha_{1},\ldots,\alpha_{L_{3}}\}\subset[0.1,10]$. We found this was fast and performed well in simulations for both continuous and binary outcomes, across a wide range of $N$ and $n_{i}$ values. We anticipate the relative speedups of using fast cluster CV, compared to standard K-fold CV, would be grow as $p$ increases (e.g. from increasing the number of knots or covariates, $q$). This is because fast K-fold inverts a $p\times p$ matrix only once per $\Lambda_{1}$ value, whereas standard K-fold inverts a similar $p\times p$ matrix for every unique $\{k,\Lambda_{1}\}$ pair.
A.7 Fast Wild Cluster Bootstrap
Motivated by classic (Cameron et al., 2008) and score-based wild cluster bootstrapping (Kline and Santos, 2010) approaches, we apply a fast wild cluster bootstrap to calculate quantiles to construct Wald 95% CIs. This can also be described as a special case of a cluster multiplier bootstrap with Rademacher weights. We describe our approach in Algorithm S3 below. Since all bootstrap replicate one-step estimates use the same initial estimate $\widehat{\bm{\theta}}^{(0)}_{{\Lambda_{1}}}$ and same inverse $\bm{H}^{-1}=\left[\frac{1}{N}\sum_{i=1}^{N}\mathbb{W}_{i}(\widehat{\bm{\theta}}^{(0)}_{{\Lambda_{0}}})+\Lambda_{1}\mathbb{S}\right]^{-1}$, this bootstrapping procedure typically takes less than a second for moderately sized $p$. This is approach of keeping the above two quantities fixed is motivated by theory for cluster bootstrapping for non-functional (unpenalized) one-step GEE (Cheng et al., 2013) as well as for score-based wild cluster bootstrapping (Kline and Santos, 2010).
[…]
Algorithm S3 Wild cluster bootstrap critical values for joint/pointwise CIs for one-step fGEE
1:Inputs:
  • Pivot coefficient estimate $\widehat{\bm{\theta}}^{(0)}_{\bm{\Lambda}_{0}}\in\mathbb{R}^{p}$ (initial estimate; used for bootstrap centering).
  • Center (target) coefficient estimate $\widehat{\bm{\theta}}^{(1)}_{\bm{\Lambda}_{1}}\in\mathbb{R}^{p}$ (one-step or final estimate to center reported bands).
  • Penalty matrix $\bm{P}\coloneqq\bm{\Lambda}_{1}\,\mathbb{S}\in\mathbb{R}^{p\times p}$.
  • Cluster-wise quantities evaluated at $\widehat{\bm{\theta}}^{(0)}_{\bm{\Lambda}_{0}}$: $\{\mathbb{W}_{i}(\widehat{\bm{\theta}}^{(0)}_{\bm{\Lambda}_{0}}),\mathbf{b}_{i}(\widehat{\bm{\theta}}^{(0)}_{\bm{\Lambda}_{0}})\}_{i=1}^{N}$, where
    • $\mathbb{W}_{i}(\widehat{\bm{\theta}}^{(0)}_{\bm{\Lambda}_{0}})\in\mathbb{R}^{p\times p}$,
    • $\mathbf{b}_{i}(\widehat{\bm{\theta}}^{(0)}_{\bm{\Lambda}_{0}})\in\mathbb{R}^{p}$.
  • Evaluation grid $\{s_{\ell}\}_{\ell=1}^{L}\subset\mathcal{S}$ and term-wise linear maps $\{\mathbb{A}_{r}\}_{r=1}^{q}$ with $\mathbb{A}_{r}\in\mathbb{R}^{L\times p}$.
  • Bootstrap replicates $B$, nominal level $\alpha\in(0,1)$.
  • Covariance estimate $\widehat{\mathrm{Var}}(\widehat{\bm{\theta}}^{(1)}_{\bm{\Lambda}_{1}})=\widehat{\bm{\Sigma}}\in\mathbb{R}^{p\times p}$ for studentization.
2:Precomputation:
[…]
5:Compute and store $\bm{H}^{-1}$.
[…]
7:Compute pivot curves, center curves, and fixed SE curves (term-wise):
8:for $r=1$ to $q$ do
9:  $\widehat{\bm{\beta}}^{\mathrm{pivot}}_{r}\leftarrow\mathbb{A}_{r}\widehat{\bm{\theta}}^{(0)}_{\bm{\Lambda}_{0}}\in\mathbb{R}^{L}$.
10:  $\widehat{\bm{\beta}}^{\mathrm{ctr}}_{r}\leftarrow\mathbb{A}_{r}\widehat{\bm{\theta}}^{(1)}_{\bm{\Lambda}_{1}}\in\mathbb{R}^{L}$.
11:  $\widehat{\mathbf{s}}_{r}\leftarrow\sqrt{\mathrm{diag}\!\big(\mathbb{A}_{r}\,\widehat{\bm{\Sigma}}\,\mathbb{A}_{r}^{T}\big)}\in\mathbb{R}^{L}$.
12:Wild cluster bootstrap: generate studentized processes
13:for $b=1$ to $B$ do
14:  Draw i.i.d. Rademacher multipliers $\{\xi_{i}^{(b)}\}_{i=1}^{N}$ with $\mathbb{P}(\xi_{i}^{(b)}=\pm 1)=1/2$.
15:  $\mathbf{B}^{(b)}\leftarrow\frac{1}{N}\sum_{i=1}^{N}\xi_{i}^{(b)}\,\mathbf{b}_{i}(\widehat{\bm{\theta}}^{(0)}_{\bm{\Lambda}_{0}})\in\mathbb{R}^{p}$.
16:  $\widehat{\bm{\theta}}^{(b)}\leftarrow\widehat{\bm{\theta}}^{(0)}_{\bm{\Lambda}_{0}}+\bm{H}^{-1}\left(\mathbf{B}^{(b)}-\mathbf{p}\right)$ (one-step bootstrap update; fixed bread)
17:  for $r=1$ to $q$ do
18:   $\bm{\beta}_{r}^{(b)}\leftarrow\mathbb{A}_{r}\widehat{\bm{\theta}}^{(b)}\in\mathbb{R}^{L}$.
19:   $\mathbf{T}_{r}^{(b)}\leftarrow\left(\bm{\beta}_{r}^{(b)}-\widehat{\bm{\beta}}^{\mathrm{pivot}}_{r}\right)\oslash\widehat{\mathbf{s}}_{r}\in\mathbb{R}^{L}.$ (studentized process; centered at pivot)
20:   $M_{r}^{(b)}\leftarrow\max_{\ell\in[L]}\left|T_{r,\ell}^{(b)}\right|$.
21:   Store pooled pointwise magnitudes $\{|T_{r,\ell}^{(b)}|:\ell\in[L]\}$ for scalar pointwise calibration.
22:Critical values (per curve $r$):
23:for $r=1$ to $q$ do
24:  $c^{\mathrm{jt}}_{r}\leftarrow\mathrm{Quantile}_{1-\alpha}\left(\{M_{r}^{(b)}\}_{b=1}^{B}\right)$.
25:  $c^{\mathrm{pt}}_{r}\leftarrow\mathrm{Quantile}_{1-\alpha}\left(\{|T_{r,\ell}^{(b)}|:\;b\in[B],\ell\in[L]\}\right)$.
26:(Optional) EDF/t-adjustment of critical values:
27:for $r=1$ to $q$ do
28:  $\mathrm{df}_{r}\leftarrow\max(\mathrm{df}_{\min},\,N-\mathrm{edf}_{r})$.
29:  $a_{r}\leftarrow\dfrac{t_{1-\alpha/2,\mathrm{df}_{r}}}{z_{1-\alpha/2}}$.
30:Construct confidence bands (reported bands centered at $\widehat{\bm{\beta}}^{\mathrm{ctr}}_{r}$):
31:for $r=1$ to $q$ do
32:  $\widehat{\beta}^{\mathrm{ctr}}_{r}(s_{\ell})\ \pm\ a_{r}c^{\mathrm{jt}}_{r}\,\widehat{s}_{r}(s_{\ell}),\qquad\ell=1,\ldots,L.$
33:  $\widehat{\beta}^{\mathrm{ctr}}_{r}(s_{\ell})\ \pm\ a_{r}c^{\mathrm{pt}}_{r}\,\widehat{s}_{r}(s_{\ell}),\qquad\ell=1,\ldots,L.$
34:return $\{a_{r}c^{\mathrm{jt}}_{r},~a_{r}c^{\mathrm{pt}}_{r}\}_{r=1}^{p}$ and the corresponding joint/pointwise scalar bands.
A.8 Fast Cluster Bootstrap Variance Estimator
@@
Motivated by theory for cluster bootstrapping in non-functional one-step GEE (Cheng et al., 2013), we propose a fast cluster bootstrap as an alternative method to estimate $\text{Var}(\widehat{\bm{\theta}}_{{\Lambda_{1}}}^{(1)})$, or to construct non-parametric bootstrap-based joint CIs. Namely, for bootstrap replicate, $t$